the base of a triangle exceeds the height by 9 yards, if the area is 221 square yards, find the length of the base and the height of the triangle
Answer by mananth(16946)

let height be h
base = h+9
area = 1/2 * h * (h+9)
221= 1/2 (h^2+9h)
442= h^2 +9h
h^2+9h-442=0
h^2+26h-17h-442=0
h(h+26)-17(h+26)=0
(h-17)(h+26)=0
So h = 17 yards the height
base = 26 yards
...
m.ananth@hotmail.ca